# SplatOverflow: Asynchronous Hardware Troubleshooting

### v2

@@ Abstract.
As tools for designing and manufacturing hardware become more accessible, smaller producers can develop and distribute novel hardware. However, there aren’t established tools to support end-user hardware troubleshooting or routine maintenance. As a result, technical support for hardware remains ad-hoc and challenging to scale. Inspired by software troubleshooting workflows like StackOverflow, we propose a workflow for asynchronous hardware troubleshooting: SplatOverflow. SplatOverflow creates a novel boundary object, the SplatOverflow scene, that users reference to communicate about hardware. The scene comprises a 3D Gaussian Splat of the user’s hardware registered onto the hardware’s CAD model. The splat captures the current state of the hardware, and the registered CAD model acts as a referential anchor for troubleshooting instructions. With SplatOverflow, maintainers can directly address issues and author instructions in the user’s workspace. The instructions define workflows that can easily be shared between users and recontextualized in new environments. In this paper, we describe the design of SplatOverflow, detail the workflows it enables, and illustrate its utility to different kinds of users. We also validate that non-experts can use SplatOverflow to troubleshoot common problems with a 3D printer in a user study.
@@ 1. Introduction
Hardware design and manufacturing tools have enabled small teams to develop and distribute novel machines and devices for niche applications. Examples of such hardware range from automation equipment for mid-scale manufacturing, such as the Lumen pick-and-place, to personal fabrication machines, such as the Prusa MK3S, to do-it-yourself gadgets like the Open Book e-reader. We refer to such products generally as hardware. For producers of this kind of hardware, turning a prototype into a viable product and supporting a growing user base with varying technical expertise remains challenging (Hodges and Chen, 2019). This paper focuses on one aspect of producing hardware: supporting end-user troubleshooting and maintenance.
Detailed and thorough documentation is an essential part of developing a hardware product(Bonvoisin et al., 2017; Ackermann, 2008; Hodges and Chen, 2019). However, documentation alone is insufficient to support the long tail of niche hardware issues users may encounter (ORR, 1996a). Instead, Subbaraman and Peek (2023a) argue that maintenance should be considered a core part of owning such hardware, and systems should be designed to support end-user troubleshooting and maintenance. We take inspiration from the infrastructure that supports software maintenance and troubleshooting workflows to examine how to create similar systems for hardware.
@@
Hardware needs a robust boundary object to support the asynchronous modes of communication necessary to scale maintenance and troubleshooting infrastructure. We present SplatOverflow, a system that enables asynchronous hardware troubleshooting through a novel boundary object: a SplatOverflow scene. A SplatOverflow scene captures the as-built hardware through a scan (a 3D Gaussian Splat (Kerbl et al., 2023)) and aligns that scan to a shared CAD model of the hardware. The scan lets a remote user independently navigate a local user’s environment and inspect issues with the as-built hardware. The CAD model lets users interact with individual parts of the hardware and manipulate them to make actions. Interactions with the SplatOverflow scene are captured in a timeline that captures the history of instructions and discussions relating to a hardware issue.
The construction of a SplatOverflow scene presents distinct advantages to local users (who have the hardware in front of them), remote assistants, and future users who may encounter similar hardware issues. Local users benefit from access to technical documentation linked to the CAD model and the ability to describe issues on their hardware by pointing and clicking on parts rather than knowing hardware-specific vocabulary. Remote assistants benefit from seeing the local user’s issue registered onto the familiar CAD model and being able to freely move the view-port to inspect the hardware in the local user’s environment. Finally, future users benefit from being able to retrieve and replay solutions to past issues without seeking out synchronous support. Instead, SplatOverflow can re-contextualize instructions in a past issue by overlaying them onto the current user’s environment. This ability to accommodate multiple users and assistants is essential to how SplatOverflow can help support scaling the maintenance effort for hardware. In this paper, we present the design of SplatOverflow, demonstrate how it can be used to troubleshoot issues on different kinds of hardware, and validate that non-expert users can use SplatOverflow to troubleshoot hardware issues.
@@ 2. Walkthrough
We demonstrate how a local user can troubleshoot a complex issue on the Lumen v3 Pick-and-Place machine using SplatOverflow. The Lumen is an automation machine that can assemble PCBs by picking up surface-mount electrical components with a vacuum nozzle and placing them on circuit boards. In this scenario, the local user’s machine can’t reliably pick up parts with one of the two vacuum nozzles. They use SplatOverflow to capture the issue they are facing and receive suggestions from a remote maintainer.
@@ 2.1. Creating a SplatOverflow Scene
They open SplatOverflow on their smartphone to capture the issue they are facing. Figure 2 shows SplatOverflow’s capture interface. The local user records a roughly one-minute-long video, filming the Lumen from multiple angles. After recording, they upload the video, and SplatOverflow generates a scene.
After two minutes, SplatOverflow completes generating a low resolution scene. This scene is generated using down-sampled video, which yields a lower-quality splat. However, this lower-resolution scene allows the user to quickly begin using SplatOverflow. After roughly four minutes from the original upload, SplatOverflow finishes generating the full-resolution scene and presents the user with a rendering of their Lumen registered onto its CAD model. SplatOverflow removes the background regions of the splat by default to preserve the local user’s privacy. What’s left is a splat of the user’s hardware and parts of the immediate work surface on which it is placed. Figure 3 shows the scene as it is rendered in the browser.
@@ 2.2. Visually Querying Documentation and Past Issues
First, the local user examines any technical documentation for Lumen to find a solution. They select SplatOverflow’s documentation tool and click on the nozzle in the scene. Figure 4(a) shows how SplatOverflow then displays sections of the technical documentation and past SplatOverflow issues referencing the nozzle assembly in the CAD model. The user scrolls through the retrieved documentation but doesn’t find any suggestions about troubleshooting poor suction.
Next, the user sees a past issue that references a faulty nozzle. They open the issue, and SplatOverflow re-projects the instructions from that scene onto their hardware. Figure 4(b) shows the solution to a prior issue reprojected into the local user’s SplatOverflow scene. The user walks through the solution, which had to do with fixing the O-rings that seal the vacuum nozzle. They check that their O-rings aren’t dislodged and confirm this is not the source of their problem. The user then decides to capture more details about the issue they are experiencing and ask for help.
@@ 2.3. Asking for Help
To share their issue with a remote maintainer, the local user adds a description of the issue they are facing. They also use SplatOverflow’s video annotation gesture to add a video of the nozzle failing to pick up a component. SplatOverflow places the video into the scene to match the filming perspective, as shown in Figure 5a. The video describes the dynamic aspects of the issue that cannot be captured in the scan of the workspace. The local user posts the issue and waits for a remote maintainer to view it and offer suggestions.
@@ 2.4. Providing Guidance
After a few hours, the remote maintainer reviews the SplatOverflow issue and offers their feedback. Figure 5b shows the maintainer reviewing the issue by and inspecting the nozzle assembly in the scene. The maintainer cannot see any issues with the nozzle assembly and suspects that the source of the issue may be the pneumatics under Lumen’s staging plate. However, the underside of the staging plate is not visible in the machine’s current orientation.
The maintainer directs the local user to reorient the machine using SplatOverflow’s transform gesture and requests that they update the captured splat after moving it. The move instruction is visualized by animating the CAD model to move from the original aligned position to a new target position. The request to update the splat includes a QR code pointing the user to SplatOverflow’s mobile capture interface for adding a splat to an existing scene. Figure 6 shows the two instructions authored by the maintainer in the timeline and how each is rendered in SplatOverflow.
@@ 2.5. Understanding Suggestions
The local user reviews the suggestions made by the remote maintainer. They see that the maintainer has asked them to move their machine. Clicking the transform event in the SplatOverflow timeline shows the local user that they must stand the machine on the back legs. Figure 6 shows the user following the maintainer’s instructions by moving the machine and updating their splat. The user scans the QR code in the timeline event and records a video capturing the underside of the staging plate. Once the splat is updated, the scene contains two splats, with the machine in different orientations. Figure 7(a) shows the updated scene with the Lumen standing upright and the pneumatics easily visible for the remote maintainer to inspect. The splat captures the state of the wiring and routed tubes that are not included in the machine’s CAD model.
@@ 2.6. Back and Forth Communication
The maintainer sees the updated splat and looks closely at the pneumatics under the staging plate. The remote maintainer inspects the underside of the pneumatics. They can use the splat to inspect the routing of tubes not modeled in the CAD and the state of components that may be degrading. After orbiting around the model, the maintainer feels confident the wiring is correct; however, the maintainer notices two things. First, the rubber bands holding the pumps have wholly degraded, as shown in 7(c). Second, the blue fitting on one of the pneumatic fittings has come loose and will likely not provide an adequate seal, as shown in 7(d). Both are potential sources of the issue, so the remote maintainer instructs the user to fix both. They use SplatOverflow’s point-to gesture and attach a comment to the pneumatic fitting as well as the vacuum pumps. They instruct the local user to test if the tubes leaving the air pump are still securely attached and to replace the rubber bands when they have the chance. Next, the maintainer guides the user on testing if air is escaping the fitting when the pump is turned. The maintainer asks the user to place their finger on the seal and feel for air when the pump is on. In case the seal leaks air, the maintainer suggests that the local user apply some hot glue as a stop-gap solution while waiting for a replacement. Figure 8 shows these gestures rendered in the timeline and within the scene.
The local user reviews this suggestion and performs the test as described by the maintainer. When they turn on the pump, they feel air escaping the fitting. They apply some hot glue, and the vacuum pressure becomes more substantial. The local user orders a replacement fitting and replies to the suggestion indicating that a poor fitting was the culprit. Once the issue is resolved, the timeline is indexed to all the parts referenced in the back-and-forth exchange. This way, if a future user has a similar issue, the timeline for how to test this fitting will be available for them to reference.
@@ 3. Utility of SplatOverflow
We contribute a technique to construct a boundary object comprising a 3D Gaussian Splat aligned to a CAD model and a timeline that can record interactions with the fused artifact (Kerbl et al., 2023). We implement a system, SplatOverflow, that uses this technique to enable asynchronous troubleshooting workflows for hardware. This allows for clear communication between parties and enables troubleshooting and maintenance workflows to be shared between users. We argue that this workflow offers unique value to local users experiencing hardware issues and remote maintainers helping fix issues.
For the local user, the SplatOverflow scene introduces two kinds of interactions: retrieving digital information indexed to the CAD model and rendering instructions, physical actions, and references to hardware parts in situ. In SplatOverflow, users can click on the part they are interested in to retrieve information about it, such as assembly documentation. To do this, SplatOverflow uses its mapping from scan points to CAD parts to identify what the user has selected and search for references to that part in the available documentation. This search method can also surface past issues that other users have troubleshooted in SplatOverflow. This form of search allows users to query for information without knowing machine-specific vocabulary. For users who aren’t familiar with the design of the hardware, this can simplify the process of navigating the corpus of documentation accompanying the hardware. SplatOverflow can also reduce the effort of mapping instructions across different contexts by rendering actions and references in the same environment as the user’s hardware. Rather than translating a step from an online guide into action on their hardware, users can view the instruction as an overlay in their SplatOverflow scene. This can help clarify instructions that require physical actions to be carried out on the hardware. Moreover, since SplatOverflow anchors interactions to the CAD model of the hardware, instructions authored in one SplatOverflow issue can be reprojected onto any scan of the same hardware. In practice, this means users can replay instructions from past issues overlaid onto their own hardware.
SplatOverflow allows remote maintainers to independently navigate another user’s hardware and rapidly author instructions. The registered 3D scan in SplatOverflow allows the remote maintainer to view the user’s hardware from novel viewpoints. This lets the maintainer diagnose the actual state of the hardware, and the aligned CAD model allows them to quickly examine how the hardware may differ from the original design. Once a maintainer is ready to make suggestions, they can use SplatOverflow’s gestures to author them directly in the scene. These gestures can communicate how the user should manipulate parts, where they should look, or what additional information the remote maintainer needs. Each gesture is attached to the SplatOverflow timeline and rendered with an accompanying visual overlaid onto the scene. By enabling these interactions, SplatOverflow allows the remote maintainer to provide support without coordinating with the end user or other maintainers.
Moreover, while we have referred to the maintainer as a singular person, in practice, multiple people could carry out this role. A SplatOverflow scene can capture the history of the user’s issue, suggestions made by multiple maintainers, and the back-and-forth about what worked and what did not. This ability to accommodate multiple users and maintainers is essential to how SplatOverflow can help support scaling the maintenance effort for hardware.

### v3

@@ Abstract.
As tools for designing and manufacturing hardware become more accessible, smaller producers can develop and distribute novel hardware. However, processes for supporting end-user hardware troubleshooting or routine maintenance aren’t well defined. As a result, providing technical support for hardware remains ad-hoc and challenging to scale. Inspired by patterns that helped scale software troubleshooting, we propose a workflow for asynchronous hardware troubleshooting: SplatOverflow.
SplatOverflow creates a novel boundary object, the SplatOverflow scene, that users reference to communicate about hardware. A scene comprises a 3D Gaussian Splat of the user’s hardware registered onto the hardware’s CAD model. The splat captures the current state of the hardware, and the registered CAD model acts as a referential anchor for troubleshooting instructions. With SplatOverflow, remote maintainers can directly address issues and author instructions in the user’s workspace. Workflows containing multiple instructions can easily be shared between users and recontextualized in new environments.
In this paper, we describe the design of SplatOverflow, the workflows it enables, and its utility to different kinds of users. We also validate that non-experts can use SplatOverflow to troubleshoot common problems with a 3D printer in a usability study.
Project Page: https://amritkwatra.com/research/splatoverflow.
Figure 1. This figure is a workflow diagram showing the different steps that go into troubleshooting with a SplatOverflow scene. It has eight components labeled (a) through (h) moving from (a) on the left to (h) on the right. At the left extreme of the figure there are four components that make up the SplatOverflow scene: A user-captured scan, existing technical documentation, a database of past issues and a CAD model of the hardware. Subfigure A depicts these elements coming together to create a SplatOverflow scene. Subfigure B shows how users can point and click on parts in the scene to retrieve their associated documentation or past issues that reference the part. Subfigure C shows how a user adds a video annotation and seeks out a maintainer’s help. Subfigure D has two parts, on the left a maintainer reviews the end-user’s issue on their screen, on the right, the SplatOverflow interface shows the suggestion authored by the maintainer. Subfigure E shows the user viewing the suggestion in situ, executing the fix, and updating the scene. Subfigure F shows the maintainer reviewing the updated scan, comparing the as-built hardware to the CAD design and proposing suggestions. The figure has a dashed line at this point, indicating that the issue could have more back-and-forth between the end-user and the maintainer. Subfigure G picks up when the end-user has identified the issue and executes the solution. Subfigure H loops back to the beginning, illustrating how resolved issues feed back into the database of past issues for future users to reference.
@@ 1. Introduction
Figure 2. This figure has three subfigures in it: A, B, C, and D. The overall figure shows the SplatOverflow interface. The left box A highlights the menu of buttons that represent gestures SplatOverflow offers for selecting parts, guiding attention, and communicating actions. The middle label B shows the 3D Gaussian Splat of the hardware with the wireframe CAD model registered and aligned with the splat. The middle-right subfigure C shows SplatOverflow’s timeline that captures the troubleshooting interaction as a sequence of instructions and responses between users. The farthest right subfigure D features a user holding a mobile phone with SplatOverflow’s mobile interface.
Hardware design tools have enabled small teams to develop and distribute novel machines and devices for niche applications. Examples of this type of hardware range from automation equipment for mid-scale manufacturing, such as the Lumen pick-and-place111https://www.opulo.io/products/lumenpnp, to personal fabrication machines, like the Prusa MK3S222https://help.prusa3d.com/tag/mk3s-2, to do-it-yourself gadgets like the Open Book333https://www.oddlyspecificobjects.com/projects/openbook/ e-reader. We refer to such products generally as hardware. For hardware producers, turning a prototype into a viable product entails addressing challenges that stem from supporting a growing user base with varying technical expertise (Hodges and Chen, 2019). This paper focuses on one aspect of producing hardware: supporting end-user troubleshooting and maintenance.
Detailed and thorough documentation is an essential part of developing a hardware product (Bonvoisin et al., 2017; Ackermann, 2008; Hodges and Chen, 2019). However, documentation alone is insufficient to support the long tail of niche hardware issues consumers may encounter (ORR, 1996). Instead, Subbaraman and Peek (2023a) argue that maintenance should be considered a core part of owning this kind of hardware, and systems should be designed to support end-user troubleshooting and maintenance. We take inspiration from the infrastructure that supports software maintenance and troubleshooting workflows to examine how to create similar systems for hardware.
@@
Hardware needs a robust boundary object to support the asynchronous modes of communication necessary to scale maintenance and troubleshooting infrastructure. We present SplatOverflow, a system that enables asynchronous hardware troubleshooting through a novel boundary object: a SplatOverflow scene. A SplatOverflow scene captures the as-built hardware through a scan (a 3D Gaussian Splat (Kerbl et al., 2023)) and renders as-designed details by aligning the scan to a shared CAD model of the hardware.
The construction of a SplatOverflow scene presents distinct advantages to local users (who have the hardware in front of them), remote maintainers, and future users who may encounter similar hardware issues. Local users benefit from access to technical documentation linked to the CAD model and the ability to describe issues on their hardware by pointing and clicking on parts rather than knowing hardware-specific vocabulary. Remote maintainers benefit from seeing the local user’s issue registered onto the familiar CAD model and being able to freely move through the scene to inspect the hardware in the local user’s environment. Finally, future users benefit from being able to retrieve and replay solutions to past issues without seeking out support. Instead, SplatOverflow re-contextualizes instructions from a past issue by overlaying them onto the current user’s environment. This ability to accommodate multiple users and maintainers is essential to how SplatOverflow can help support scaling the maintenance effort for hardware.
In this paper, we present the design of SplatOverflow, demonstrate how it can be used to troubleshoot issues on different kinds of hardware, and validate that non-expert users can use SplatOverflow to troubleshoot hardware issues.
@@ 2. Walkthrough
We demonstrate how a non-expert local end-user can troubleshoot a complex issue on the Lumen v3 Pick-and-Place machine using SplatOverflow. The Lumen is an automation machine designed to assemble PCBs by picking up surface-mount electrical components with a vacuum nozzle and placing them on circuit boards. In this scenario, the local user’s machine can not reliably pick up parts with one of the two vacuum nozzles. They use SplatOverflow to capture the issue they are facing and receive suggestions from a remote maintainer. Figure 1 depicts a high-level visual illustration of this process.
@@ 2.1. Creating a SplatOverflow Scene
The user opens SplatOverflow on their smartphone to capture the issue they are facing. Figure 2(d) shows SplatOverflow’s capture interface. They record a roughly one-minute-long video, filming the Lumen from multiple angles. After recording, they upload the video, and SplatOverflow generates a scene.
After two minutes, SplatOverflow completes generating a low-resolution scene. This scene is generated using down-sampled video, which yields a lower quality splat. However, this lower-resolution scene allows the user to quickly begin using SplatOverflow. After roughly four minutes from the original upload, SplatOverflow finishes generating the full-resolution scene containing the user’s Lumen registered onto its CAD model. SplatOverflow removes the background regions of the splat by default to preserve the local user’s privacy. What’s left is a splat of the user’s hardware and parts of the immediate work surface on which it is placed. Figure 2(a-c) shows the scene as it is rendered in the browser.
Figure 3. This figure has two subfigures in it: A and B. The left subfigure shows an example of the SplatOverflow documentation query interface. It features a user clicking on the nozzle of the machine in SplatOverflow and the corresponding nozzle documentation being displayed on the right side of the screen. The right subfigure shows an example of asynchronous collaboration using SplatOverflow’s timeline feature. With the user selecting the nozzle of the pick and place machine. Various past issues other users have had are displayed on the right side of the interface. The figure also features two arrow annotations pointing towards the nozzle being overlaid on the user’s machine, with the arrows being from a past user’s suggestions.
@@ 2.2. Visually Querying Documentation and Past Issues
First, the local user examines any technical documentation for Lumen to find a solution. They select SplatOverflow’s documentation tool and click on the nozzle in the scene. Figure 3(a) shows how SplatOverflow then displays sections of the technical documentation and past SplatOverflow issues referencing the nozzle assembly. The user scrolls through the retrieved documentation but does not find any suggestions about troubleshooting poor suction.
Next, the user sees a past issue that references a faulty nozzle. They open the issue, and SplatOverflow re-contextualizes the instructions from that scene onto their hardware. Figure 3(b) shows the solution to a past issue recontextualized in the local user’s SplatOverflow scene. The user walks through the solution, which had to do with fixing the O-rings that seal the vacuum nozzle. They check that their O-rings are not dislodged and confirm this is not the source of their problem. They then decide to capture more details about the issue and ask for help.
Figure 4. This figure has two subfigures in it: A and B. The left subfigure A shows a local user’s video annotation rendered in the SplatOverflow scene of a pick and place machine. The video is placed as a floating screen in the 3D Scene and aligned to the same perspective it was filmed in. The right subfigure B shows the remote maintainer inspecting the vacuum nozzle in the SplatOverflow scene.
@@ 2.3. Asking for Help
To share their issue with a remote maintainer, the local user adds a description of the problem they are facing. They also add a video annotation of the nozzle failing to pick up a component. SplatOverflow places the video into the scene to match the filming perspective, as shown in Figure 4(a). The video describes the dynamic aspects of the issue that cannot be captured in the scan of the workspace. The local user then posts the issue and waits for a remote maintainer to view it and offer suggestions.
@@ 2.4. Providing Guidance
After a few hours, the remote maintainer reviews the SplatOverflow issue and offers feedback. Figure 4(b) shows the maintainer reviewing the issue and inspecting the nozzle assembly in the scene. The maintainer cannot see any problems with the nozzle assembly and suspects that the source of the issue may be the pneumatics under Lumen’s staging plate. However, the underside of the staging plate is not visible in the machine’s current orientation.
The maintainer directs the local user to reorient the machine by manipulating the CAD model in SplatOverflow and requests that they update the captured splat after moving it. The move instruction is visualized by animating the CAD model to move from the original position to a new target position. The request to update the splat includes a QR code pointing the user to SplatOverflow’s mobile capture interface for adding a splat to an existing scene. Figure 5 shows the instructions authored by the maintainer in the timeline and how each is rendered in SplatOverflow.
Figure 5. This figure has four subfigures in it: A, B, C, and D. The left subfigure A shows the wireframe corresponding to the final position a pick and place machine should be in for maintenance. The middle subfigure B shows a green button in the SplatOverflow timeline that describes the issue and also plays a CAD animation on how to move the machine once clicked. The other middle subfigure C is a green popup in the timeline that offers a QR code for the local user to capture a new splat of a machine in the new orientation described in B. The right subfigure D shows a local user performing the action specified by the maintainer and preparing to update the splat.
Figure 6. This figure has four subfigures in it: A, B, C, and D. The middle subfigure A shows an update of the previous splat with the pick and place machine now in the orientation specified by the remote maintainer. The right subfigure B shows that the SplatOverflow timeline has been updated to indicate that a new splat has been added. The upper left subfigure C shows a close up that highlights a part in the splat that the remote maintainer is concerned about, the vacuum pumps on the underside of the machine. The lower left subfigure D shows a close up that highlights aother part in the splat that the remote maintainer is concerned about a pneumatic fitting.
@@ 2.5. Understanding Suggestions
The local user reviews the suggestions made by the remote maintainer. They see that the maintainer has asked them to move their machine. Clicking the instructions in the SplatOverflow timeline shows the user that they need to stand the machine on its back legs. Figure 5(d) shows the user following the maintainer’s instructions by moving the machine and updating their splat. The user scans the QR code in the timeline event and captures a new splat showing the underside of the staging plate. Once the splat is updated, the scene contains two splats, with the machine in different orientations. Figure 6(a) shows the updated scene with the Lumen standing upright and the pneumatics easily visible for the remote maintainer to inspect. The splat captures the state of the wiring and routed tubes that are not included in the machine’s CAD model.
Figure 7. This figure has four subfigures in it: A, B, C, and D. Subfigure A is a green arrow pointing towards a part that the remote maintainer references. Subfigure B is a box in the timeline with the relevant instruction for the part highlighted in Subfigure A. Subfigure C shows the CAD wireframe overlaid onto missing parts in the Splat with Subfigure D giving relevant instructions on fixes.
@@ 2.6. Back and Forth Communication
The maintainer sees the updated splat and examines the pneumatics under the staging plate. Using the splat, they inspect the routing of tubes (not modeled in the CAD) and assess the condition of components that may be degrading. After orbiting around the model, the maintainer feels confident that the wiring is correct but notices two concerns. First, the rubber bands holding the pumps have wholly degraded, as shown in Figure 6(c). Second, the blue fitting on one of the pneumatic fittings has come loose and will likely not provide an adequate seal, as shown in Figure 6(d). Both are potential sources of the issue, so the remote maintainer instructs the user to fix both. They use SplatOverflow’s pointing gesture to attach comments to the pneumatic fitting and the vacuum pumps. They instruct the local user to test if the tubes leaving the air pump are still securely attached and to replace the rubber bands when they have the chance. Next, the maintainer guides the user through testing if air escapes the fitting when the pump is turned on. The maintainer asks the user to place their finger on the seal and feel for air when the pump is on. In case the seal leaks air, the maintainer suggests that the local user apply some hot glue as a stop-gap solution while waiting for a replacement part. Figure 7 shows how these instructions are rendered in the timeline and within the scene.
The local user reviews the suggestion and performs the test as described by the maintainer, as shown in Figure 8. When they turn on the pump, they feel air escaping the fitting. They apply some hot glue, and the vacuum pressure becomes more substantial. The local user orders a replacement fitting and replies to the suggestion indicating that a poor fitting was the culprit. Once the issue is resolved, the timeline is indexed to all the parts referenced in the back-and-forth exchange.
Figure 8. This figure has three subfigures in it: A, B, and C. The overall figure shows a local user following the asynchronous instructions of a remote maintainer. The right subfigure A shows the local use testing whether the tubing on the machine is still firmly attached to the vacuum pump. The middle subfigure B shows the local user testing whether they can feel any air escaping the fitting that the remote maintainer referenced. The right subfigure C shows the local user trying the stop-gap solution of applying some hot glue to the fitting to improve the seal.
@@ 3. Utility of SplatOverflow
This section outlines the capabilities made possible by SplatOverflow. We discuss the scenarios in which SplatOverflow addresses the limitations of existing troubleshooting artifacts, highlight the novel capabilities SplatOverflow enables, and outline the utility it creates for different users.
3.1. Existing Troubleshooting Artifacts
Existing hardware troubleshooting workflows use a variety of artifacts to facilitate back-and-forth communication between parties, each with its own drawbacks. We highlight how SplatOverflow mitigates these drawbacks by complementing each kind of artifact with additional information and context within a scene.
3.1.1. Text Posts
Text is commonly used to describe the physical state of hardware, how it may be erroring, or what action must be taken on the hardware. However, text descriptions can create confusion if parties do not share a common vocabulary. This is especially true when non-expert users are troubleshooting hardware with maintainers. Non-expert users may not know how to refer to specific parts or which part a provided instruction refers to. SplatOverflow explicitly links text to the referent parts from the CAD model and visually highlights corresponding parts when the text is selected. This preserves the flexibility of text while addressing issues of referential uncertainty and ambiguity.
3.1.2. Annotated Images
Annotated images direct attention to a specific area or part of the hardware. Moreover, annotations allow maintainers to specify instructions through deictic references on the local user’s hardware, e.g., tighten this bolt (where this bolt is circled in an image). This lets local users reason about instructions without changing contexts, as the annotation is authored by referencing their hardware. However, image annotations can only reference what the local user captured in the image. When the image does not capture the appropriate details, maintainers must guide the user without context cues from the local user’s hardware. SplatOverflow extends the utility of image annotation for maintainers by allowing them to annotate the scene from various arbitrary viewpoints.
3.1.3. Video Tutorials
Video tutorials can capture detailed multi-step processes on hardware and show local users how to manipulate their hardware through demonstration. Video tutorials contain a large amount of detail, but can demand significant time and planning to be effective. As a result, video documentation is often used for build instructions or set-up guides, the utility of which can be pre-empted and planned for. However, troubleshooting is often about problems that are hard to preempt (ORR, 1996). Creating detailed video tutorials for each of these issues as they arise is demanding for maintainers. In contrast, impromptu and unedited videos can be difficult for users to parse, leading to more miscommunication. SplatOverflow improves the utility of impromptu video filmed by non-expert users by situating the filming perspective of the video within the scene.
3.2. Novel Capabilities
In addition to addressing some of the limitations of existing artifacts, SplatOverflow enables novel capabilities that are not supported by existing asynchronous hardware troubleshooting artifacts.
3.2.1. Indexing Issues
In SplatOverflow, resolved issues are indexed in a database based on the parts they reference. Future users can subsequently query this database by selecting the parts they are experiencing issues with. This streamlines accessing the evolving technical information related to the hardware. Similar to the pools of technical knowledge created by the social network of copier repair technicians described by Orr (ORR, 1996), the indexed corpus of hardware issues created by SplatOverflow is a valuable reference for other maintainers and future users (Bobrow and Whalen, 2002).
3.2.2. Independent Asynchronous Navigation
With SplatOverflow, maintainers can independently navigate and author instructions from perspectives and on parts not intentionally captured by the local user. This adds flexibility and enables maintainers to inspect and investigate the actual state of the local user’s hardware.
3.2.3. Issue Re-contextualization
By defining annotations in reference to CAD data, SplatOverflow can re-contextualize them into new scenes containing the same hardware. This allows local users to view suggestions from past issues overlaid onto their own hardware and reduces the amount of context shifting needed to make sense of instructions from past issues.
3.3. Utility to Users
SplatOverflow mediates an interaction between two kinds of users, non-expert local end-users and expert remote maintainers. This section underscores the utility to each type of user and the conditions in which SplatOverflow is most beneficial.
3.3.1. For End-Users
SplatOverflow provides the most utility to users whose expertise is not in the hardware itself but in what they use it for. These users depend on the hardware for productivity but are not familiar with the intricacies of its design. As such, maintenance and troubleshooting are simultaneously critical and challenging tasks for these users to carry out. SplatOverflow benefits these end-users in three ways: (1) By indexing past issues and relevant technical information, SplatOverflow helps end-users search for workflows to remedy issues they are facing. (2) By re-contextualizing past issues, SplatOverflow allows end-users to reason about and follow existing instructions within the context of their own hardware. (3) By capturing the hardware context in a scan, SplatOverflow allows end-users to seek help without knowing hardware-specific terminology. These benefits are accessed by capturing a brief video of their hardware.
3.3.2. For Maintainers
As hardware develops, end-user maintenance and troubleshooting is essential to support (Dunn et al., 2023; ORR, 1996; Hodges and Chen, 2019; Eghbal, 2020). This entails hardware designers, producers, or enthusiasts taking on the role of maintainers to support the growing number of non-expert end-users. SplatOverflow helps maintainers scale support for hardware in three ways. (1) By creating an asynchronous troubleshooting workflow, SplatOverflow allows maintainers to address end-user issues at their own pace. (2) By registering an immersive scan to a CAD model, SplatOverflow allows maintainers to compare the actual state of the hardware to its design without context shifting across artifacts. (3) By indexing resolved issues, SplatOverflow allows maintainers to inspect and share links to related issues and relevant solutions.
These features help maintainers attend to individual issues. However, as more issues are addressed and indexed through SplatOverflow, the benefits to maintainers expand. For example, the maintainer’s effort in troubleshooting issues is multiplied when future end-users can reference the solution without seeking out help. Moreover, when multiple people are responsible for maintenance and troubleshooting, they can reference the pool of advice offered by others in the past while diagnosing new issues. As such, SplatOverflow supports maintainers in scaling hardware support.